Lindenauer Corp. bought a machine on January 1, 2008 for $800,000. The machine had an expected life of 20 years and was expected to have a salvage value of $40,000. The company does not plan to dispose of the machine but does believe it may be impaired. On July 1, 2018, the company reviewed the potential of the machine and determined that its future net cash flows totaled $350,000 and its fair value was $230,000.

Calculate the impairment loss, if any, as of July 1, 2018 and prepare any journal entry needed
Answer #1

Machine purchased on January 01, 2008

Purchase price = $ 800,000 , Useful Life = 20 years

Salvage Value = $ 40,000 , Depreciable value = 800,000 - 40,000 = $ 760,000

Depreciation per year = $ 760,000 / 20 = $ 38,000

On July 01, 2008, the company reviewed the potential of the machine :

period between Jan 01, 2008 and July 01, 2018. = 10.5 years

Book Value of machine (July 01, 2008) = $ 800,000 - ( $38,000 X 10.5) = $ 800,000 - $ 399,000

Book value of machine (July 01, 2008) = $ 401,000.

Impairment is nothing but the difference between Book value of Asset and the Amount recoverable from the

Asset.

Amount recoverable from the Machine = Higher of Future net Cash Flows and Its Fair Value

Amount recoverable from the Machine = Higher of $ 350,000 and $ 230,000 = $ 350,000

Impairment Loss = Book value of Asset - the Amount recoverable from the Machine

Impairment Loss as of July 01, 2018 = $ 401,000 - $ 350,000 = $ 51,000

Journal Entry

Impairment Loss a/c Dr.... $ 51,000

To Machinery $ 51,000